OpenStudy (anonymous):

find the total ares of all shaded rectangles L=1 and W=a. I started like this 1a + 1a=(1a)^2, (1 +a)(1+1)(1+a)(a+a) is this on the right track? If not show me how to do this please

OpenStudy (anonymous):

the shaded areas of the rectangles length 1 and width a

OpenStudy (anonymous):

We know that a rectangle has an area of length * width 

OpenStudy (anonymous):

The problem states that that the length is 1 and that the width is "a" Using the equation 1a+1a=1a^2 we can determine what a equal \[1a+1a=1a^2\] \[2a=a^2\] \[2=a\] Hence we can conclude that a=2 Therefore our area of the triangle which is length* width =1*2=2